# DocHPLT: A Massively Multilingual Document-Level Translation Dataset

### v1

@@ 4 Experiments and Findings
To demonstrate the viability of our dataset, we first identify the optimal context length for fine-tuning on DocHPLT across varying approaches in existing DocMT research. We then compare this best-performing fine-tuned configuration against off-the-shelf instruction-tuned models on the same test set. Finally, we investigate monolingual and multilingual fine-tuning for DocMT on DocHPLT. In the following sections, the notation of src-trg refers to the src-to-trg translation direction.
@@
We test English translation into a total of 10 languages, for their diversity in script, typology, and resource availability, as well as their inclusion in WMT24++ (Deutsch et al., 2025) (for testing) and CometKiwi (Rei et al., 2022) (for filtering). The languages are Arabic (ar), Catalan (ca), Hindi (hi), Estonian (et), Persian (fa), Finnish (fi), Icelandic (is), Korean (kr), Malayalam (ml), and Urdu (ur). It is worth noting that generating non-English is usually harder for LLMs compared to generating English.
@@
We preprocess the documents for training by removing those with an AD below 0.3 or a doc-averaged Bicleaner score below 0.3. To further filter the documents, we apply CometKiwi (Rei et al., 2022) using SLIDE (Raunak et al., 2023) so that a candidate translation’s context compatibility can be considered by the quality estimator. We score our documents using a window of three and a slide of one, retaining those with a CometKiwi score in the top 25th percentile for every language. The aim is to ensure that only the highest quality parallel documents are used for training and evaluation.
@@
We perform supervised fine-tuning (SFT) on Qwen2.5-7B-Instruct (Qwen et al., 2025) and Llama-3.1-8B-Instruct (Grattafiori et al., 2024) using LoRA with rank 16 and alpha 32 (Hu et al., 2022), using the open-instruct toolkit. Our models are trained on 1000 documents per language, unless stated otherwise. Our hyperparameters are listed in Appendix B. At test time, we always translate an entire source document in a single pass. LLM’s chat template is always applied. All prompt information is detailed in Appendix C.
@@
We conduct evaluations on two test sets: a held-out set from DocHPLT and WMT24++, selected for their overlapping language coverage. We construct DocHPLT test by randomly sampling 500 documents per language from the CometKiwi-filtered corpora. We rigorously de-contaminate on the English side by computing Jaccard similarity over bigrams and removing any test document with a similarity above 0.8 to any training document. This ensures that our evaluation is not biased towards training on similar documents. The final test sizes are shown in Table 1.
@@
We compute BLEU444nrefs:1—case:mixed—eff:no—tok:13a—smooth:exp—version:2.5.1 (Papineni et al., 2002) and chrF++555nrefs:1—case:mixed—eff:yes—nc:6—nw:2—space:no—version:2.5.1 (Popović, 2017) scores by treating each hypothesis document and reference document as single strings, and then averaging these scores across all documents. Compared to sentence-level metrics or neural metrics, our choice of metrics 1) does not require sentence-level alignments between the hypothesis and reference documents (which is not guaranteed by the DocMT system outputs); and 2) is not limited by the context size of neural metrics.
@@ 4.1 How much context do document-level models need?
Existing research on DocMT with LLMs adopts distinct strategies to process data. Some approaches operate on a sentence level but use previous translations as context (Wu et al., 2024), some work on fixed-size chunks (Alabi et al., 2025; Wicks et al., 2024; Post and Junczys-Dowmunt, 2024), translating each chunk separately, and some perform full document-to-document translation (Ramos et al., 2025). We therefore start our experiments by training models using varying context lengths to find the best configuration.
@@
We train models with five context configurations: sentence-level (chunk 1, no context), chunks of 2, 5, and 10 sentences, as well as full document-to-document (doc2doc) training. All configurations use the same 1000 training documents. For chunk-based training, we compute the loss only on target segments while providing source context. The total number of tokens is kept constant for all languages, despite the different data formats.
@@ 4.2 Does fine-tuning on DocHPLT help document-level translation?
One key indicator of the usefulness of a data resource is whether practitioners can create better models using it. Although the origin of our data is web crawls, which may have been consumed by LLM pre-training, the parallelism signals are new in DocHPLT and not accessible through pre-training. Thus, in this section, we investigate whether we can fine-tune (FT) LLMs to achieve better DocMT results beyond prompting baselines.
@@
We fine-tune separate monolingual models for 5 languages separately: Finnish, Icelandic, Korean, Malayalam, and Urdu, using Qwen2.5-7B-Instruct and Llama-3.1-8B-Instruct on 1,000 documents per language. We use our best-performing data configuration of chunk size 10 for our experiments. Evaluation is done on held-out DocHPLT test sets and WMT24++.
@@ Results
Overall, our results suggest that off-the-shelf instruction-tuned models may already contain knowledge for these medium to low-resourced languages, yet DocMT training is still much needed to attain more desirable performance. This underscores the value of our DocHPLT, which is the first to cater to those languages in this task.
Training Languages DocHPLT WMT24++
@@ 4.3 Does multilingual training improve over monolingual models?
While our monolingual fine-tuned models achieve significant gains over prompting baselines, deploying and maintaining separate models for each language presents scalability drawbacks. Furthermore, multilingual LLM fine-tuning may offer performance advantages over monolingual tuning (Chen et al., 2024). To test whether our massively multilingual DocHPLT can be used for cross-lingual transfer in training, in this section, we build and assess multilingual models. Particularly, we test unseen languages to determine whether DocMT training’s utility extends beyond training languages.
@@ 5 Conclusion
We introduced a pipeline to derive a document-level corpus with rich metadata and presented the outcome, DocHPLT, the largest publicly available document-level translation dataset with 124 million aligned document pairs across 50 languages paired with English. The utility of our massively multilingual dataset has been demonstrated through experiments: fine-tuning LLMs on our data improved over prompting baselines, and multilingual training surpassed monolingual models, though zero-shot transfer to unseen languages remained challenging. Our experiments also revealed challenges in DocMT: full document-to-document training and generalization to other document domains. Future work may use DocHPLT data for further investigations such as large-scale training, data synthesis, and metric study.
@@ Acknowledgements
figures/hplt.pngThis project has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No 101070350 and from UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee [grant number 10052546].
@@ Appendix A Dataset Statistics
#sentences #docs
af 16,416,841 297,636
ar 65,482,300 2,271,167
az 12,202,189 332,742
be 10,672,952 212,121
bg 80,018,549 1,746,301
bn 10,473,372 414,099
bs 20,635,243 514,615
ca 47,905,003 1,198,217
cy 8,908,119 265,261
en 2,665,945,834 47,484,349
eo 6,115,355 119,196
et 33,684,509 774,561
eu 6,783,654 189,347
fa 24,837,952 810,029
fi 111,615,913 2,445,791
ga 6,398,081 172,167
gl 10,657,570 233,545
gu 3,202,679 108,507
he 38,077,820 1,190,198
hi 37,592,475 1,336,090
hr 52,267,826 1,063,347
is 12,571,982 274,078
ja 164,136,152 4,032,689
$k$ 5,948,866 140,082
kn 4,463,262 123,053
ko 84,527,642 2,058,811
lt 48,692,264 1,031,628
lv 37,426,957 796,659
mk 12,465,228 307,055
ml 2,925,457 115,189
mr 3,066,703 128,808
ms 51,150,528 978,185
mt 6,328,544 141,088
nb 89,189,502 1,884,362
ne 1,549,852 74,579
$n$ 4,228,079 93,285
si 1,497,375 50,605
sk 70,057,465 1,461,804
sl 37,501,647 797,858
sq 11,475,561 328,651
sr 21,620,629 407,440
sw 8,409,824 185,287
ta 6,790,864 215,564
te 5,131,680 141,279
th 16,134,265 676,699
tr 100,380,235 3,884,137
uk 89,841,883 1,955,041
ur 5,479,098 234,708
uz 3,502,356 69,440
vi 87,511,126 1,986,258
xh 995,556 21,561
total 4,264,894,818 87,775,169
@@ C.1 Overview
We use the same prompt for SFT and during inference for both off-the-shelf instruction-tuned and fine-tuned models.
@@ C.2 Chunk-based translation
Translate the following source segment from English into Catalan. English: [SOURCE TEXT] Catalan: [TARGET TEXT]
@@ C.3 Document-to-document translation
Translate the following source document from English into Catalan. English: [SOURCE DOCUMENT] Catalan: [TARGET DOCUMENT]

### v2

@@ 4 Experiments and Findings
In order to test the usefulness of our dataset, we apply it to the task of fine-tuning LLMs for MT. Our first set of experiments tests different context lengths for this fine-tuning to see how much performance is affected by using the larger document contexts that DocHPLT enables. We then compare this best-performing fine-tuned configuration against off-the-shelf instruction-tuned models on the same test set. Finally, we investigate monolingual and multilingual fine-tuning for DocMT on DocHPLT. In the following sections, the notation of “src-trg” refers to the src-to-trg translation direction.
@@
We test translation from English into a total of 10 languages, chosen for their diversity in script, typology, and resource availability, as well as their inclusion in WMT24++ (Deutsch et al., 2025) (for testing) and CometKiwi (Rei et al., 2022) (for filtering). The languages are Arabic (ar), Catalan (ca), Hindi (hi), Estonian (et), Persian (fa), Finnish (fi), Icelandic (is), Korean (kr), Malayalam (ml), and Urdu (ur). It is worth noting that generating non-English is usually harder for LLMs compared to generating English.
@@
We preprocess the documents for training by removing those with an AD below 0.3 or a document-averaged Bicleaner score below 0.3. We then discard unaligned segments in either source or target, and merge segments in a one-to-many alignment into a single segment. Finally, we filter data using CometKiwi (Rei et al., 2022) with SLIDE (Raunak et al., 2023): a window of 3 and a slide of 1. We retain document pairs with a CometKiwi score in the top 25th percentile for every language. This is to ensure that only high-quality parallel documents are used for training or evaluation.
@@
We perform supervised fine-tuning (SFT) on Qwen2.5-7B-Instruct (Qwen et al., 2025) and Llama-3.1-8B-Instruct (Grattafiori et al., 2024) with LoRA, rank 16 and alpha 32 (Hu et al., 2022), using the open-instruct toolkit [^4] https://github.com/allenai/open-instruct. Unless stated otherwise, our models are fine-tuned on 1000 documents per language, due to compute constraints. Our hyperparameters are listed in Appendix B. At test time, we always translate an entire source document in a single pass. LLM’s chat template is always applied. All prompt information is detailed in Appendix C.
@@
We conduct evaluations on two test sets: a held-out set from DocHPLT and WMT24++, selected for their overlapping language coverage. We construct DocHPLT test by randomly sampling 500 documents per language from the CometKiwi-filtered corpora. We de-contaminate on the English side by computing Jaccard similarity over bigrams and removing any test document with a similarity above 0.8 to any training document. This ensures that our evaluation is not biased towards training on similar documents. The final test sizes are shown in Table 1.
@@
We compute BLEU555nrefs:1—case:mixed—eff:no—tok:13a—smooth:exp—version:2.5.1 (Papineni et al., 2002) and chrF++666nrefs:1—case:mixed—eff:yes—nc:6—nw:2—space:no—version:2.5.1 (Popović, 2017) by treating each hypothesis document and reference document as a single string, and then averaging these scores across all documents. Our metric choice avoids the need for sentence-level alignment, which DocMT outputs do not guarantee. We note that while neural metrics such as COMET or LLM-as-a-judge are generally more reliable at the sentence level, their effectiveness in our document-level setting remains uncertain due to limited empirical validation, context support, and language coverage.
@@ 4.1 How much context do document-level models need?
Existing research on DocMT with LLM adopts distinct strategies to process data. Some approaches operate on a sentence level but use previous translations as context (Wu et al., 2024), some methods process fixed-size chunks (Alabi et al., 2025; Wicks et al., 2024; Post and Junczys-Dowmunt, 2024), translating each chunk separately, and some perform full document-to-document translation (Ramos et al., 2025). We start our experiments by training models using varying context lengths to find the best configuration.
@@
We build en-xx models for five target languages separately: Finnish, Icelandic, Korean, Malayalam, and Urdu. We fine-tune two open-source LLMs: Qwen2.5-7B-Instruct and Llama-3.1-8B-Instruct. We fine-tune each model under five different context configurations: sentence-level (chunk 1, no context), chunks of 2, 5, and 10 sentences, as well as full document-to-document (doc2doc) training. For chunk-based training, we compute the loss only on target segments while providing source context. The total number of tokens is kept constant for all languages, despite the different data formats.
@@ 4.2 Does fine-tuning on DocHPLT help document-level translation?
One key indicator of the usefulness of a data resource is whether practitioners can create better models using it. Although the origin of our data is web crawls, which may have been consumed by LLM pre-training, the parallelism signals are new in DocHPLT and not accessible through pre-training. Thus, in this section, we compare the results of fine-tuning LLMs to prompting baselines.
@@
We compare our fine-tuned models with the best-performing data configuration of chunk size 10 to their corresponding off-the-shelf instruction-tuned models. Evaluation is done on held-out DocHPLT test sets and WMT24++.
@@ Results
Our results suggest that off-the-shelf instruction-tuned models may already contain knowledge for these medium to low-resourced languages, yet fine-tuning on DocHPLT consistently improves performance across these languages. This underscores the value of our DocHPLT, which is the first to cater to those languages in this task. Nonetheless, we note that a higher performance does not necessarily indicate higher data quality—it may also be a result of greater exposure to a given language or to document-level input and output. A causal analysis will be useful, but for most of the languages we study, there is no suitable alternative data to compare to at the moment.
Seen Languages DocHPLT WMT24++
@@ 4.3 Does multilingual training improve over monolingual models?
While our monolingual fine-tuned models achieve significant gains over prompting baselines, deploying and maintaining separate models for each language presents scalability drawbacks. Furthermore, multilingual LLM fine-tuning may offer performance advantages over monolingual tuning (Chen et al., 2024). To test whether our multilingual DocHPLT can be exploited for cross-lingual transfer in training, in this section, we build and assess multilingual models. Particularly, we test on both seen and unseen languages to determine whether benefits extend beyond training languages.
@@ 5 Conclusion
We introduced a pipeline to derive a document-level corpus with rich metadata and presented the outcome, DocHPLT, the largest publicly available document-level translation dataset with 124 million aligned document pairs across 50 languages paired with English. The utility of our massively multilingual dataset has been demonstrated through experiments: fine-tuning LLMs on our data improved over prompting baselines, and multilingual training surpassed monolingual models, though zero-shot transfer to unseen languages remained challenging. Our experiments also revealed challenges in DocMT: full document-to-document training and generalization to other document domains. Future work may use DocHPLT data for further investigations such as large-scale training, data filtering, data synthesis, and DocMT metric study.
@@ Acknowledgements
figures/eu-logo.pngThis project has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No 101070350 and from UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee [grant number 10052546].
Dayyán O’Brien is also supported by a G-Research NextGen Scholarship, part of the UKRI AI Centre for Doctoral Training in Responsible and Trustworthy in-the-world Natural Language Processing (grant ref: EP/Y030656/1).
@@ Appendix A Dataset Statistics
#sentences #docs #deduped docs
af 16,416,841 297,636 286,861
ar 65,482,300 2,271,167 2,196,334
az 12,202,189 332,742 321,500
be 10,672,952 212,121 203,758
bg 80,018,549 1,746,301 1,669,696
bn 10,473,372 414,099 405,339
bs 20,635,243 514,615 488,093
ca 47,905,003 1,198,217 1,131,468
cy 8,908,119 265,261 253,040
en 2,665,945,834 47,484,349 45,995,228
eo 6,115,355 119,196 103,858
et 33,684,509 774,561 747,075
eu 6,783,654 189,347 175,318
fa 24,837,952 810,029 785,963
fi 111,615,913 2,445,791 2,341,993
ga 6,398,081 172,167 166,516
gl 10,657,570 233,545 215,009
gu 3,202,679 108,507 106,423
he 38,077,820 1,190,198 1,149,349
hi 37,592,475 1,336,090 1,315,174
hr 52,267,826 1,063,347 1,009,227
is 12,571,982 274,078 265,088
ja 164,136,152 4,032,689 3,934,457
$k$ 5,948,866 140,082 135,689
kn 4,463,262 123,053 120,996
ko 84,527,642 2,058,811 2,003,338
lt 48,692,264 1,031,628 995,288
lv 37,426,957 796,659 766,138
mk 12,465,228 307,055 292,992
ml 2,925,457 115,189 111,721
mr 3,066,703 128,808 126,552
ms 51,150,528 978,185 942,418
mt 6,328,544 141,088 137,104
nb 89,189,502 1,884,362 1,809,266
ne 1,549,852 74,579 73,691
$n$ 4,228,079 93,285 78,426
si 1,497,375 50,605 48,730
sk 70,057,465 1,461,804 1,406,726
sl 37,501,647 797,858 765,171
sq 11,475,561 328,651 317,315
sr 21,620,629 407,440 386,829
sw 8,409,824 185,287 178,185
ta 6,790,864 215,564 208,573
te 5,131,680 141,279 138,727
th 16,134,265 676,699 655,628
tr 100,380,235 3,884,137 3,767,266
uk 89,841,883 1,955,041 1,891,287
ur 5,479,098 234,708 228,952
uz 3,502,356 69,440 68,191
vi 87,511,126 1,986,258 1,940,095
xh 995,556 21,561 20,797
total 4,264,894,818 87,775,169 84,882,858
@@ C.1 Overview
We use the same prompt for SFT and during inference for both off-the-shelf instruction-tuned and fine-tuned models. LLM’s chat template is always applied.
@@ C.2 Chunk-based translation
Translate the following source segment from [SOURCE LANGUAGE] into [TARGET LANGUAGE]. [SOURCE LANGUAGE]: [SOURCE TEXT] [TARGET LANGUAGE]: [TARGET TEXT]
@@ C.3 Document-to-document translation
Translate the following source document from [SOURCE LANGUAGE] into [TARGET LANGUAGE]. [SOURCE LANGUAGE]: [SOURCE DOCUMENT] [TARGET LANGUAGE]: [TARGET DOCUMENT]